Direction: Solve the following problems stated below. Show your solutions.



1. Compute the electric flux of the following sets of variables:


a. E = 7.2x10^-5. N/C


A = 1.6x10^-6 m^2


θ = 30°



b. E = 9.7x10^5 N/C


A = 5.3x10^-6 m^2


θ = 75°



c. E = 6.5x10^5 N/C


A = 8.3x10^-6 m^2


θ = 30°


1
Expert's answer
2022-02-02T05:15:31-0500

1. Compute the electric flux of the following sets of variables:

a. E = 7.2x10^5. N/C

A = 1.6x10^-6 m^2

θ = 30°


"\phi=EA\cos\theta\\
{\rm=7.2*10^{5}\:N/C*1.6*10^{-6}\: m^2\cos30^\circ}\\
=1.0\;\rm N\cdot m^2/C"


b. E = 9.7x10^5 N/C

A = 5.3x10^-6 m^2

θ = 75°



"\phi=EA\cos\theta\\
{\rm=9.7*10^{5}\:N/C*5.3*10^{-6}\: m^2\cos75^\circ}\\
=1.3\;\rm N\cdot m^2/C"



c. E = 6.5x10^5 N/C

A = 8.3x10^-6 m^2

θ = 30°


"\phi=EA\cos\theta\\
{\rm=6.5*10^{5}\:N/C*8.3*10^{-6}\: m^2\cos30^\circ}\\
=4.7\;\rm N\cdot m^2/C"